OpenStudy (anonymous):

If the problem type is labeled “ClickAndPlace”, click the item to select it and then click the location where it belongs; otherwise, click the item and drag it into place. Express answer in exact form. Find the area of the smaller segment whose chord is 8" long in a circle with an 8" radius GIVING OUT MEDALS IF ANYONE HELPS ME !!!!

OpenStudy (anonymous):

Chord AB = 8 radius OA = 8 draw perpendicular sinø = AM/OA = 4/8 = 1/2 ø = 30° <AOB = 2 × 30 = 60° area of arc = 60/360 × π × 8² = 33.49 for area of triangle OAB OM = √(8² – 4²) = 6.93 area of triangle = 1/2 × 6.93 × 8 = 27.71 area of minor segment = 33.49 – 27.71 = 5.78 in²